The no. Of ways in which the six faces of a cube be painted with six different colours?

Grade:12

2 Answers

Supraja
6 years ago
Each side can be painted in 6 different ways with six different colours.No. Of ways = 6x6x.....6times=6^6=46,656ways
Supraja
6 years ago
Sorry for the wrong answer above, Each side can be painted in six different ways with six different colours. No. Of ways=6+6+6+6+6+6=36